• When a box of mass 45kg is given as the initial speed pf 5m/s, it slides along a horizontal floor distance of 3m before coming to rest. What is the coefficient of dynamic friction between the box and the floor.
1
Expert's answer
2021-10-27T10:16:58-0400

"F_f=ma\to \mu mg=m(v^2/2s)\to\mu=v^2/(2sg)=5^2/(2\cdot3\cdot9.8)=0.43"